# Active learning for regression in engineering populations: A risk-informed approach

### v1

@@ Abstract
Regression is a fundamental prediction task common in data-centric engineering applications that involves learning mappings between continuous variables. In many engineering applications (e.g. structural health monitoring), feature-label pairs used to learn such mappings are of limited availability which hinders the effectiveness of traditional supervised machine learning approaches. The current paper proposes a methodology for overcoming the issue of data scarcity by combining active learning with hierarchical Bayesian modelling.
Active learning is an approach for preferentially acquiring feature-label pairs in a resource-efficient manner. In particular, the current work adopts a risk-informed approach that leverages contextual information associated with regression-based engineering decision-making tasks (e.g. inspection and maintenance). Hierarchical Bayesian modelling allow multiple related regression tasks to be learned over a population, capturing local and global effects. The information sharing facilitated by this modelling approach means that information acquired for one engineering system can improve predictive performance across the population.
The proposed methodology is demonstrated using an experimental case study. Specifically, multiple regressions are performed over a population of machining tools, where the quantity of interest is the surface roughness of the workpieces. An inspection and maintenance decision process is defined using these regression tasks which is in turn used to construct the active-learning algorithm. The novel methodology proposed is benchmarked against an uninformed approach to label acquisition and independent modelling of the regression tasks. It is shown that the proposed approach has superior performance in terms of expected cost – maintaining predictive performance while reducing the number of inspections required.
keywords:
active learning; regression; risk assessment; structural health monitoring; population modelling
articletype: RESEARCH ARTICLEjournal: Data-Centric Engineering
[…]
{Frontmatter}
[1]Daniel R. Clarkson
[…]
Daniel R. Clarkson et al.
[…]
{policy}
[Impact Statement]
This paper addresses online learning of regression models in a cost-effective manner. The problem is addressed via two avenues; active learning; i.e. choosing information critical to decisions, and information sharing, using a hierarchical Bayesian model. This paper presents a decision-theoretic approach that allows the monetary benefits to be quantified exactly. Regression is a prominent and important tool in many decision-support technologies such as health monitoring and digital twins. Improving performance when learning regressions in an online manner could lead to reduced costs and improved safety in many science and engineering applications.
@@ 1 Introduction
Statistical pattern recognition (SPR) is widely recognised as the primary tool for data-driven predictions in SHM systems Farrar and Worden, (2012). Regression models are a fundamental component of the SPR approach to decision-support technologies, enabling the prediction of continuous outcomes based on acquired data. For example, within the context of SHM, by associating the target variables of a regression model with salient health-states, inferences can be made regarding the condition of a structure of interest - although assumed as discrete in many SHM problems, damage progression is usually, in reality, a continuous variable.
A significant challenge in SHM is the scarcity of data. All data-driven models can suffer from bias and high uncertainty without sufficient training data, leading to unreliable predictions. Acquiring extensive labelled datasets capturing a structure’s behavior across various health conditions is prohibitively costly and often unattainable for essential infrastructures. Inadequate data motivate sharing information between similar assets Bull et al., (2023). From this a new approach has emerged, Population-based structural health monitoring (PBSHM) Bull et al., (2021); Gosliga et al., (2021); Gardner et al., (2021); Tsialiamanis et al., (2021).
@@ 1.1 Population-based Structural Health Monitoring
PBSHM considers an entire population of structures. This approach assumes that structures that share common environmental conditions, load patterns, and/or aging effects and thus will share statistical commonalities. Considering structures as a population allows SHM users to share data between them. It allows a structure with rich historical data to lend its statistical strength to a data-poor structure. By capturing data from multiple structures, PBSHM aims to provide a more accurate and holistic health assessment and enables the identification of global patterns, trends and anomalies that would be difficult to observe at the individual level. This strategy can not only improve predictions for structures with very little data, but also makes the most out of comprehensive datasets that can be so costly to procure. PBSHM has a suite of tools that can be used to share information across a population, namely transfer learning and domain adaptation. For a thorough introduction to the population-based approach to SHM see Bull et al., (2021); Gosliga et al., (2021); Gardner et al., (2021); Tsialiamanis et al., (2021).
@@ 1.2 Transfer Learning
In PBSHM, one of the most widely-explored areas of transfer learning is known as domain adaptation in which feature data are mapped from a label-rich source domain to a label-scarce target domain, with the aim of reducing the distance between the source and target domain in a shared latent space such that label-information can be shared. Fink et al. discuss domain adaptation’s place in fleet prognostics and health management Fink et al., (2020). Gardner et al. utilised domain adaptation to transfer inferences across different structures, considering a population of laboratory-scale three-story buildings Gardner et al., (2020). Zhang et al. used domain adaptation for fault diagnosis in the context of rotating machinery and between different machines Zhang et al., (2017); Li et al., (2020). Xu et al. showed that transfer learning can be used to diagnose story-wise damage conditions of buildings effected by earthquakes Xu and Noh, (2021).
@@ 1.3 Active Learning
Active learning (AL) presents another avenue to address the limitations imposed by data scarcity in SHM. Conventional supervised machine-learning methods are infeasible for many SHM applications because of the costs associated with descriptive labels. This has led to the development of unsupervised and semi-supervised machine-learning techniques. Active learning is a set of techniques that selectively queries labels for otherwise unlabelled data that are most informative given the current model; the model can then be updated using this informed subset of labelled data. Active learning can be applied offline to a large pool of collected data Wang et al., (2017), or online, whereby the dataset is continuously updated as new observations are collected Zhu et al., (2007). The online setting is particularly significant in SHM; generally data from a monitored structure will become available gradually throughout the life of the structure. In many cases, inspecting monitored systems can be extremely costly, so if a system can determine when only the most critical or informative observations need to be investigated, this could lead to significant reductions in maintenance costs Bull et al., 2019b .
Active learning has seen growing interest, particularly within SHM; Bull et al. provide an online active learning framework for the classification problem and show the effects of the framework via a case study on acoustic emissions data Bull et al., 2019b . Hughes et al. present a risk-based formulation of active learning in which queries are guided by the expected value of information Hughes et al., 2022a and outline a method to minimise the effects of sampling bias in active learning Hughes et al., 2022b .
In the literature, most research on active learning focusses on the classification setting. However, for many real-life applications, such as degradation Shahraki et al., (2017) and crack growth He et al., (2023), regression-based models are more suitable. The work done in the area of active learning for regression (ALR) is comparatively much less. One of the first statistical analyses of ALR was provided by Cohn et al., (1996), where a locally-weighted regression model for studying the dynamics of a robot arm is showcased. More recently, Wu et al., (2019) proposed a ’greedy sampling’ method for active learning regression. In that paper a greedy sampling method proposed by Yu and Kim, (2010), is expanded to be ’greedy’ in the output space and is shown to be a robust and effective method for active learning. Cai et al., (2013) proposed an expected model change maximisation framework for regression, the learner chooses the unlabelled instance which causes the maximum change in the current model parameter. Freund et al., (1997) showed that query-by-committee is not only applicable to binary labels but also to discrete labels. For more thorough surveys of active learning the reader is directed to Kumar and Gupta, (2020); Fu et al., (2013); Aggarwal et al., (2014).
@@ 2.1 Hierarchical Bayesian Modelling
Section 2.1 follows an explanation provided by the authors precious work Bull et al., (2023). Consider data recorded from a population of K engineering structures. The population can be denoted,
@@
and the output $y_{ik}$ is determined by evaluating one of $K$ latent functions with seperate additive noise $\epsilon_{k}$. The mapping $f_{k}$ is assumed to be correlated between members in the population. The models should be improved by learning the parameters in a joint inference over the whole population. In machine learning this is referred to as multi-task learning (MTL); in statistics, such data are usually modelled with hierarchical models Gelman and Hill, (2006); Kreft and De Leeuw, (1998).
In practice, some members in a population may possess extensive historical data, while members that may have been more recently deployed will have very limited data for training. Learning separate independent models for each group might lead to unreliable results for data-poor members, while a single regression model of all the data (complete pooling) would result in poor generalisation. Hierarchical Bayesian models can learn separate models for each member, while encouraging the parameters of these models to be correlated Murphy, (2012). The established theory is summarised here.
@@
The parent nodes $\left\{\mathbf{\mu}_{\alpha},{\mathbf{\sigma}^{2}}_{\alpha}\right\}$ are inferred from the data so Equations (6)–(8) encode prior belief of the dependence between latent variables. If these parent nodes were a fixed value, rather than inferred, each model would be independent, preventing the flow of information between domains, this structure allows data-sparse domains to borrow statistical strength from those that are data rich Murphy, (2012).
Huang and Beck, (2015); Huang et al., (2019) present an early example of using Hierarchical Bayesian models to represent engineering structures for SHM. Bull et al., (2023) used hierarchical models to improve the survival analysis of a truck fleet and power prediction in a wind farm. Di Francesco et al., (2021) used hierarchical models to account for incomplete and imperfect data in an inspection-planning setting.
@@ 2.2 Decision Theory
The idea of a ’rational’ decision-maker is defined as a decision-maker that acts to maximise the expected-utility of their actions; this is expressed via the von Neumann-Morgenstern theorem Morgenstern et al., (1964), which states that, for two decidable actions $a$ and $b$:
@@
where $\succeq$ denotes a weak preference, indicating that a decision-maker favours $a$ at least as much $b$; and $EU(\cdot)$ denotes the expected utility associated with doing an action. In Morgenstern et al., (1964), the expected utility is derived. Consider a stochastic event $X$. Which has mutually-exclusive outcomes of which $x\in\mathcal{X}$ are conditionally dependent on a decision $D$ between actions $a$ and $b$. The expected utility of action $a$ is computed as follows,
@@ 2.3 Active learning
To implement a Bayesian hierarchical model, such as the one proposed in Section 2.1, labelled data is required. As discussed in Section 1, for many engineering applications, and particularly for SHM, collecting labels for data is very expensive. It requires sending a domain expert to inspect the physical asset and often requires the operation of the asset to be temporarily halted. These costs motivate reducing the number of inspections and only inspecting measurements that would most improve the predictive model. Active learning is a set of tools that do this. Additionally, an offline active learner would not be suitable for SHM. SHM systems are required to analyse data as it arrives throughout the life of the monitored system and decisions about interventions can be required immediately upon arrival of this data. One of the major challenges of an online active learner is that once the decision is made not to query a label, access is lost to this information. It is not possible to retrieve the label once the opportunity has passed.
Consider data, $X={\left\{\mathbf{x}_{i}\right\}^{N}}_{i=1}$, which have hidden labels, $Y={\left\{\mathbf{y}_{i}\right\}^{N}}_{i=1}$, which can be acquired by paying the costs associated with inspection. The process of choosing and labelling these data points is referred to as querying. An active learner aims to learn a mapping of the observations, $X$, to the labels, $Y$, while keeping queries to a minimum. A general heuristic for active learning is presented by Bull et al., 2019a and adapted to regression below,
Figure 2: An active learning heuristic adapted from Bull et al., 2019a
The measure that an active learner uses to decide which unlabelled data to query is important. Information theoretic approaches use information measures such as entropy and uncertainty to guide querying, these types of approaches are common. Hughes et al., 2022a suggest a risk-based active learner for SHM. This approach uses the expected utility of an action, or the ’Value of Information’, to guide querying. While risk-based active learning has been explored in the classification setting Hughes et al., 2022b , the current paper proposes an approach to risk-based active learning for the regression problem in which queries are guided according to the expected utility.
@@ 3.1 The dataset
A dataset described in the authors previous work Wickramarachchi, (2019) measures deterioration over the life of machining tools during a turning process. The experimental set-up is shown in Figure 3. The workpiece is rotated around the dashed A-line and the tool makes four passes along the workpiece. Each pass starts at point S and ends at point E. After four passes, the tool is inspected, and measurements are taken of the workpiece and tool. This process continued until tool failure. The investigation was repeated for seven nominally identical tools, which form a population.
Figure 3: Schematic showing the experimental set-up used for data acquisition Wickramarachchi, (2019)
@@ 3.2 The hierarchical model
Because of the natural degradation of tools during the machining process, and its effect on surface finish, tools must be replaced regularly. While each tool may be produced to the same specification and made from the same materials, there will be variations in the manufacturing process that manifest as variations between the physical properties of the tools and differences in behaviour between the tools; this can be an issue for standard modelling techniques. However, this variation lends itself well to a hierarchical Bayesian model because these types of model account for variations within a population while taking advantage of the statistical similarities between them. An additional benefit of hierarchical models is their suitability to the online setting and sparse datasets. This is particularly useful for tool condition monitoring where researchers may need to make predictions as soon as the machining process has begun and with only a few data points to learn a model. Additionally, the usual benefits of Bayesian modelling apply (uncertainty quantification, prior information etc.). The hierarchical framework set out in Section 2.1 is adapted below to the case study. Here, there is a population of $K=7$ similar tools, with a target response vector $y_{k}$, the roughness measurements for each tool. The input vectors $x_{k}$ are the sliding distance measurements for each tool (how far the tool has cut across the work piece). $\{x_{ik},y_{ik}\}$ are the $i^{\text{th}}$ pair of observations in tool $k$. There are $N_{k}$ observations in each member and thus $\Sigma^{K}_{k=1}N_{k}$ observations in total. The aim is to learn a set of $K$ predictors related to the regression task. The type of hierarchical model used for this analysis is a linear mixed model Kreft and De Leeuw, (1998), so for each member in the population, a gradient $m_{k}$ and intercept $c_{k}$ are learnt. A graphical representation of this model can be seen in Figure 5.
$y_{ik}$$x_{ik}$$m_{k}$$c_{k}$$\gamma_{k}$$\mu_{m}$$\sigma_{m}$$\mu_{c}$$\sigma_{c}$$\gamma$$k$$\theta$$s_{\sigma_{m}}$$s_{\sigma_{c}}$$\bar{\mu}_{c}$$s_{\mu_{c}}$$s_{\gamma}$$i\in(1,..,N)$$k\in(1,..,K)$
@@
Where the slopes are $\mathrm{normally}$ distibuted, with mean ${\mu}_{m}$ and standard deviation $\sigma_{m}$. Equation 17 shows the prior expectation of the slopes is $\mathrm{Gamma}$ distributed with shape $k=1$ and scale $\theta=1$. The Gamma distribution was a natural choice because we want to encode that the gradients are always positive over the life of a tool. Equation 18 shows that the prior deviation of the slope is $\mathrm{HalfCauchy}$ distributed with location parameter equal to zero and scale parameter $s_{\sigma_{m}}=25$. As recommended by Gelman, (2006), the variance priors for this hierarchical model are set to be weakly informative. The priors for the intercepts are
@@
The systematic application of graph-theoretic algorithms has led to a number of probabilistic programming languages. Here, models are implemented in NumPyro Phan et al., (2019). The parameters are inferred using MCMC, via the no U-turn implementation of Hamiltonian Monte Carlo Hoffman and Gelman, (2014). Throughout, the burn-in period is 1000 iterations and 2000 iterations are used for inference.
To motivate a population-based approach, the hierarchical model will be compared to a model with complete-pooling and an independent model with no pooling at all. In the complete-pooling approach, a single regression for all the data is learnt. For the independent model, a regression is learnt for each member in the population but it is assumed there is no correlation between them. Hierarchical modelling is somewhere in between, where a separate regression can be learnt for each member, while encouraging the parameters of these models to be correlated Murphy, (2012).
@@
Table 1: A table containing the mean squared error of partial-pooling, complete-pooling and no-pooling methods
Mean Squared Error Pooling Type Tool 1 Tool 2 Tool 3 Tool 4 Tool 5 Tool 6 Tool 7 Total No-Pooling 0.03539 0.00848 0.02781 0.00541 0.00930 0.08680 0.06081 0.23401 Complete-Pooling 0.03844 0.01492 0.03578 0.00716 0.00506 0.06062 0.02525 0.18724 Partial-Pooling 0.03928 0.00629 0.03167 0.00523 0.00575 0.06099 0.02677 0.17599
These results are in accordance with the results of the authors previous work Bull et al., (2023); Dardeno et al., (2024) and motivates sharing information across a population to improve predictions.

### v2

@@ Abstract
Regression is a fundamental prediction task common in data-centric engineering applications that involves learning mappings between continuous variables. In many engineering applications (e.g. structural health monitoring), feature-label pairs used to learn such mappings are of limited availability which hinders the effectiveness of traditional supervised machine learning approaches. The current paper proposes a methodology for overcoming the issue of data scarcity by combining active learning with hierarchical Bayesian modelling. Active learning is an approach for preferentially acquiring feature-label pairs in a resource-efficient manner. In particular, the current work adopts a risk-informed approach that leverages contextual information associated with regression-based engineering decision-making tasks (e.g. inspection and maintenance). Hierarchical Bayesian modelling allow multiple related regression tasks to be learned over a population, capturing local and global effects. The information sharing facilitated by this modelling approach means that information acquired for one engineering system can improve predictive performance across the population. The proposed methodology is demonstrated using an experimental case study. Specifically, multiple regressions are performed over a population of machining tools, where the quantity of interest is the surface roughness of the workpieces. An inspection and maintenance decision process is defined using these regression tasks which is in turn used to construct the active-learning algorithm. The novel methodology proposed is benchmarked against an uninformed approach to label acquisition and independent modelling of the regression tasks. It is shown that the proposed approach has superior performance in terms of expected cost – maintaining predictive performance while reducing the number of inspections required.
Keywords: active learning; regression; risk assessment; structural health monitoring; population modelling
@@ 1   Introduction
Statistical pattern recognition (SPR) is widely recognised as the primary tool for data-driven predictions in SHM systems [10]. Regression models are a fundamental component of the SPR approach to decision-support technologies, enabling the prediction of continuous outcomes based on acquired data. For example, within the context of SHM, by associating the target variables of a regression model with salient health-states, inferences can be made regarding the condition of a structure of interest - although assumed as discrete in many SHM problems, damage progression is usually, in reality, a continuous variable.
A significant challenge in SHM is the scarcity of data. All data-driven models can suffer from bias and high uncertainty without sufficient training data, leading to unreliable predictions. Acquiring extensive labelled datasets capturing a structure’s behavior across various health conditions is prohibitively costly and often unattainable for essential infrastructures. Inadequate data motivate sharing information between similar assets [2]. From this a new approach has emerged, Population-based structural health monitoring (PBSHM) [3, 18, 14, 33].
@@ 1.1   Population-based Structural Health Monitoring
PBSHM considers an entire population of structures. This approach assumes that structures that share common environmental conditions, load patterns, and/or aging effects and thus will share statistical commonalities. Considering structures as a population allows SHM users to share data between them. It allows a structure with rich historical data to lend its statistical strength to a data-poor structure. By capturing data from multiple structures, PBSHM aims to provide a more accurate and holistic health assessment and enables the identification of global patterns, trends and anomalies that would be difficult to observe at the individual level. This strategy can not only improve predictions for structures with very little data, but also makes the most out of comprehensive datasets that can be so costly to procure. PBSHM has a suite of tools that can be used to share information across a population, namely transfer learning and domain adaptation. For a thorough introduction to the population-based approach to SHM see [3, 18, 14, 33].
@@ 1.2   Transfer Learning
In PBSHM, one of the most widely-explored areas of transfer learning is known as domain adaptation in which feature data are mapped from a label-rich source domain to a label-scarce target domain, with the aim of reducing the distance between the source and target domain in a shared latent space such that label-information can be shared. Fink et al. discuss domain adaptation’s place in fleet prognostics and health management [11]. Gardner et al. utilised domain adaptation to transfer inferences across different structures, considering a population of laboratory-scale three-story buildings [15]. Zhang et al. used domain adaptation for fault diagnosis in the context of rotating machinery and between different machines [39, 27]. Xu et al. showed that transfer learning can be used to diagnose story-wise damage conditions of buildings effected by earthquakes [37].
@@ 1.3   Active Learning
Active learning (AL) presents another avenue to address the limitations imposed by data scarcity in SHM. Conventional supervised machine-learning methods are infeasible for many SHM applications because of the costs associated with descriptive labels. This has led to the development of unsupervised and semi-supervised machine-learning techniques. Active learning is a set of techniques that selectively queries labels for otherwise unlabelled data that are most informative given the current model; the model can then be updated using this informed subset of labelled data. Active learning can be applied offline to a large pool of collected data [34], or online, whereby the dataset is continuously updated as new observations are collected [40]. The online setting is particularly significant in SHM; generally data from a monitored structure will become available gradually throughout the life of the structure. In many cases, inspecting monitored systems can be extremely costly, so if a system can determine when only the most critical or informative observations need to be investigated, this could lead to significant reductions in maintenance costs [5].
Active learning has seen growing interest, particularly within SHM; Bull et al. provide an online active learning framework for the classification problem and show the effects of the framework via a case study on acoustic emissions data [5]. Hughes et al. present a risk-based formulation of active learning in which queries are guided by the expected value of information [23] and outline a method to minimise the effects of sampling bias in active learning [24].
In the literature, most research on active learning focusses on the classification setting. However, for many real-life applications, such as degradation [31] and crack growth [19], regression-based models are more suitable. The work done in the area of active learning for regression (ALR) is comparatively much less. One of the first statistical analyses of ALR was provided by [7], where a locally-weighted regression model for studying the dynamics of a robot arm is showcased. More recently, [36] proposed a ’greedy sampling’ method for active learning regression. In that paper a greedy sampling method proposed by [38], is expanded to be ’greedy’ in the output space and is shown to be a robust and effective method for active learning. [6] proposed an expected model change maximisation framework for regression, the learner chooses the unlabelled instance which causes the maximum change in the current model parameter. [12] showed that query-by-committee is not only applicable to binary labels but also to discrete labels. For more thorough surveys of active learning the reader is directed to [26, 13, 1].
@@ 2.1   Hierarchical Bayesian Modelling
Section 2.1 follows an explanation provided by the authors precious work [2]. Consider data recorded from a population of K engineering structures. The population can be denoted,
@@
and the output $y_{ik}$ is determined by evaluating one of $K$ latent functions with seperate additive noise $\epsilon_{k}$. The mapping $f_{k}$ is assumed to be correlated between members in the population. The models should be improved by learning the parameters in a joint inference over the whole population. In machine learning this is referred to as multi-task learning (MTL); in statistics, such data are usually modelled with hierarchical models [17, 25].
In practice, some members in a population may possess extensive historical data, while members that may have been more recently deployed will have very limited data for training. Learning separate independent models for each group might lead to unreliable results for data-poor members, while a single regression model of all the data (complete pooling) would result in poor generalisation. Hierarchical Bayesian models can learn separate models for each member, while encouraging the parameters of these models to be correlated [29]. The established theory is summarised here.
@@
The parent nodes $\left\{\mathbf{\mu}_{\alpha},{\mathbf{\sigma}^{2}}_{\alpha}\right\}$ are inferred from the data so Equations (6)–(8) encode prior belief of the dependence between latent variables. If these parent nodes were a fixed value, rather than inferred, each model would be independent, preventing the flow of information between domains, this structure allows data-sparse domains to borrow statistical strength from those that are data rich [29].
[21, 22] present an early example of using Hierarchical Bayesian models to represent engineering structures for SHM. [2] used hierarchical models to improve the survival analysis of a truck fleet and power prediction in a wind farm. [9] used hierarchical models to account for incomplete and imperfect data in an inspection-planning setting.
@@ 2.2   Decision Theory
The idea of a ’rational’ decision-maker is defined as a decision-maker that acts to maximise the expected-utility of their actions; this is expressed via the von Neumann-Morgenstern theorem [28], which states that, for two decidable actions $a$ and $b$:
@@
where $\succeq$ denotes a weak preference, indicating that a decision-maker favours $a$ at least as much $b$; and $EU(\cdot)$ denotes the expected utility associated with doing an action. In [28], the expected utility is derived. Consider a stochastic event $X$. Which has mutually-exclusive outcomes of which $x\in\mathcal{X}$ are conditionally dependent on a decision $D$ between actions $a$ and $b$. The expected utility of action $a$ is computed as follows,
@@ 2.3   Active learning
To implement a Bayesian hierarchical model, such as the one proposed in Section 2.1, labelled data is required. As discussed in Section LABEL:sec:1, for many engineering applications, and particularly for SHM, collecting labels for data is very expensive. It requires sending a domain expert to inspect the physical asset and often requires the operation of the asset to be temporarily halted. These costs motivate reducing the number of inspections and only inspecting measurements that would most improve the predictive model. Active learning is a set of tools that do this. Additionally, an offline active learner would not be suitable for SHM. SHM systems are required to analyse data as it arrives throughout the life of the monitored system and decisions about interventions can be required immediately upon arrival of this data. One of the major challenges of an online active learner is that once the decision is made not to query a label, access is lost to this information. It is not possible to retrieve the label once the opportunity has passed.
Consider data, $X={\left\{\mathbf{x}_{i}\right\}^{N}}_{i=1}$, which have hidden labels, $Y={\left\{\mathbf{y}_{i}\right\}^{N}}_{i=1}$, which can be acquired by paying the costs associated with inspection. The process of choosing and labelling these data points is referred to as querying. An active learner aims to learn a mapping of the observations, $X$, to the labels, $Y$, while keeping queries to a minimum. A general heuristic for active learning is presented by [4] and adapted to regression below,
Figure 2: An active learning heuristic adapted from [4]
The measure that an active learner uses to decide which unlabelled data to query is important. Information theoretic approaches use information measures such as entropy and uncertainty to guide querying, these types of approaches are common. [23] suggest a risk-based active learner for SHM. This approach uses the expected utility of an action, or the ’Value of Information’, to guide querying. While risk-based active learning has been explored in the classification setting [24], the current paper proposes an approach to risk-based active learning for the regression problem in which queries are guided according to the expected utility.
@@ 3.1   The dataset
A dataset described in the authors previous work [35] measures deterioration over the life of machining tools during a turning process. The experimental set-up is shown in Figure 3. The workpiece is rotated around the dashed A-line and the tool makes four passes along the workpiece. Each pass starts at point S and ends at point E. After four passes, the tool is inspected, and measurements are taken of the workpiece and tool. This process continued until tool failure. The investigation was repeated for seven nominally identical tools, which form a population.
Figure 3: Schematic showing the experimental set-up used for data acquisition [35]
@@ 3.2   The hierarchical model
Because of the natural degradation of tools during the machining process, and its effect on surface finish, tools must be replaced regularly. While each tool may be produced to the same specification and made from the same materials, there will be variations in the manufacturing process that manifest as variations between the physical properties of the tools and differences in behaviour between the tools; this can be an issue for standard modelling techniques. However, this variation lends itself well to a hierarchical Bayesian model because these types of model account for variations within a population while taking advantage of the statistical similarities between them. An additional benefit of hierarchical models is their suitability to the online setting and sparse datasets. This is particularly useful for tool condition monitoring where researchers may need to make predictions as soon as the machining process has begun and with only a few data points to learn a model. Additionally, the usual benefits of Bayesian modelling apply (uncertainty quantification, prior information etc.). The hierarchical framework set out in Section 2.1 is adapted below to the case study. Here, there is a population of $K=7$ similar tools, with a target response vector $y_{k}$, the roughness measurements for each tool. The input vectors $x_{k}$ are the sliding distance measurements for each tool (how far the tool has cut across the work piece). $\{x_{ik},y_{ik}\}$ are the $i^{\text{th}}$ pair of observations in tool $k$. There are $N_{k}$ observations in each member and thus $\Sigma^{K}_{k=1}N_{k}$ observations in total. The aim is to learn a set of $K$ predictors related to the regression task. The type of hierarchical model used for this analysis is a linear mixed model [25], so for each member in the population, a gradient $m_{k}$ and intercept $c_{k}$ are learnt. A graphical representation of this model can be seen in Figure 5.
@@
Where the slopes are $\mathrm{normally}$ distibuted, with mean ${\mu}_{m}$ and standard deviation $\sigma_{m}$. Equation 17 shows the prior expectation of the slopes is $\mathrm{Gamma}$ distributed with shape $k=1$ and scale $\theta=1$. The Gamma distribution was a natural choice because we want to encode that the gradients are always positive over the life of a tool. Equation 18 shows that the prior deviation of the slope is $\mathrm{HalfCauchy}$ distributed with location parameter equal to zero and scale parameter $s_{\sigma_{m}}=25$. As recommended by [16], the variance priors for this hierarchical model are set to be weakly informative. The priors for the intercepts are
@@
The systematic application of graph-theoretic algorithms has led to a number of probabilistic programming languages. Here, models are implemented in NumPyro [30]. The parameters are inferred using MCMC, via the no U-turn implementation of Hamiltonian Monte Carlo [20]. Throughout, the burn-in period is 1000 iterations and 2000 iterations are used for inference.
To motivate a population-based approach, the hierarchical model will be compared to a model with complete-pooling and an independent model with no pooling at all. In the complete-pooling approach, a single regression for all the data is learnt. For the independent model, a regression is learnt for each member in the population but it is assumed there is no correlation between them. Hierarchical modelling is somewhere in between, where a separate regression can be learnt for each member, while encouraging the parameters of these models to be correlated [29].
@@
These results are in accordance with the results of the authors previous work [2, 8] and motivates sharing information across a population to improve predictions.